OpenStudy (anonymous):

Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (anonymous):

The simpliest way to do this is just to add them... Doing so would leave you with 4y (the 2x and -2x cancel out) = 20 Solve for y, and then you can solve for x, this method involves matrices.

OpenStudy (anonymous):

huuh

OpenStudy (anonymous):

See how they're lined up? y + x = number?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Adding the 3y and y give 4y, Adding 2x and -2x gives 0, and 11 and 9 make 20, simplyfing into one equation. You're basically adding one whole equation to the other.

OpenStudy (anonymous):

what did you get because I think I got something way wrong

OpenStudy (shane_b):

3y − 2x = 11 + y + 2x = 9 ----------- 4y + 0x = 20 Simplify and solve for y: 4y=20 y = 20/4 y=5 Using the second equation to solve for x (now that you know y): y + 2x = 9 5 + 2x = 9 2x = 9 - 5 2x = 4 x=4/2 x = 2